AFWEX Status, 08 December 2000

Another very good case for AFWEX.

The DC-8 had a take-off at ~1840 CST. The flight included a tight spiral ascent around the CART site to 39kft, two sets of with- and against-wind legs at ~40, 35, 31, 28, and 25kft, two sets of cross wind legs at the same altitudes, and a tight spiral descent around the CART site. After this flight, the DC-8 has ~8.5 science flight hours remaining, and another 3 reserved for the return ferry flight. During the level legs, small angle yaw maneuvers were performed (constant yaw and compensating roll) for COAST studies.
Overall, this was a very clear night, with some interesting aerosols as well. Some variability in flight level water vapor and ozone concentrations was observed on the cross wind legs at 28 and 31 kft, due to mixing from the upper level jet. On the spiral descent, the tropopause was observed at ~33 kft, and moist layers and temperature inversions at ~33 and ~28 kft.
The Proteus flew its last AFWEX mission today, with a ~1845 CST take-off, followed by a slow (500 ft/min) ascent over CART site, past the DC-8 to ~54 kft. A mapping pattern was performed over the CART site and the DC-8 until ~2240 CST when it headed west to overfly the western boundary facility and then continued on its ferry flight back to California. A very successful deployment for the Proteus !

At the 3pm science meeting, preparation for the upcoming ARM Science Team Meeting in March and AFWEX data submission and distribution was discussed briefly. Shortly after the experiment ends, an email with details will be sent out to the AFWEX participants.

At midday, expecting possibility of mid and high clouds and also some low thick clouds. By early evening, mid and upper level clouds near the Kansas/Oklahoma border were evaporating and moving east; mostly clear skies were found across northern Oklahoma. By ~1645 CST, severe clear conditions were observed visually over CART site. At ~1900 (just after take-off time), an isolated thick cloud formed at ~6km only over the CART site but quickly dissipated and moved east (see a loop of the 3.9um GOES loop at this time). Clear conditions were found over CART site again by ~1945 CST, and persisted throughout the flights.

Instrument Status/Comments
CART site:
   CART Raman Lidar Normal operation.
   GSFC Raman Lidar Normal operation from ~1800 to 0200 CST.
   MPI DIAL Normal operation from ~1800 to 0200 CST.
   WFF Chilled Mirror Sondes Launches at 1815, 2115, and 0015 CST. 9 sensors remaining.
   Other Vaisala sondes launched at 1830, 2130, 0030, ... CST etc.
   LASE Normal operation.
   DLH Normal operation.
   Cryo frost-point/chilled-mirror Normal operation.
   SHIS Normal operation.
   COAST Normal operation.
   AeroSapient Normal operation.
   NAST-I Normal operation.
   NAST-M Normal operation.
   FIRSC Normal operation.

Dave Tobin, University of Wisconsin